### v2

# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

Abstract
Multimodal document understanding is a challenging task to process and comprehend large amounts of textual and visual information. Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task. However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially for academic papers. In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents. PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM. The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model. To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of English and Chinese academic papers. Multiple strategies are proposed to automatically generate 1.1 million QA pairs along with their corresponding evidence sources. Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal document understanding, surpassing proprietary products by an average of 8.6% on F1. Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
footnotetext: Equal contribution; Corresponding author
1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of PDF document understanding [1, 2], where these models have demonstrated impressive capabilities in processing and generating human-like text. However, they still face many challenges when it comes to lengthy PDF documents with interlaced text and images, such as academic papers.
To handle lengthy documents, current research in multimodal document understanding with LLMs primarily follows two mainstream technical routes. The first route is based on pure text modality understanding. As shown in Fig. 1(a), these approaches typically consider the parsed OCR results from PDF documents, and convert all visual elements into textual representations (e.g., captions and OCR content extracted from figures and tables). They then employ long-context LLMs [3, 4, 5] or utilize Retrieval Augmented Generation (RAG) techniques [6, 7, 8, 9] to process the textual content. The main disadvantage of this approach is the significant loss of visual information inherent in multimodal documents, making it challenging to support answering fine-grained visual-related questions.
Figure 1: Method comparison for long multi-page PDF document understanding. (a) Plain text solution: long-context/RAG LLMs for parsed pure text content. (b) Pure vision solution: VDU models for page-level encoding and feature interaction. (c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
The second route is based on pure visual modality understanding. As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks[10, 11, 12]. However, high-resolution images and numerous pages generate a large number of visual tokens, significantly increasing the models’ input token length and leading to scalability issues. This makes it challenging to efficiently process multi-page long documents. While some recent multi-page VDU models handle up to 8 pages [13] or 20 pages [14], they typically encode each page separately and perform page-level visual feature interactions such as concatenation [15, 16]. Nevertheless, as the number of pages grows, computational resource consumption still escalates substantially, rendering these models inefficient for processing longer documents.
Considering the limitations of existing methods in multimodal understanding of long PDF documents, we propose a new MLLM architecture with end-to-end sparse sampling, named PDF-WuKong. Since most user queries are only related to a small part of the content in a long document, the sparse sampling can significantly remove redundant noise information. It encodes text paragraphs and diagrams in the parsed PDF document, utilizing both text and image representations to identify and extract the most relevant evidence in response to a user’s query. The sampled sparse evidence significantly reduces the number of input tokens of LLM and this process is independent of the length of the input documents. Moreover, the sparse sampler and LLM can be integrated in an end-to-end manner for training and inference, optimizing the performance of multimodal representation and question answering while improving time efficiency. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs. Another important characteristic is that it can naturally provide strong interpretability for the question answering.
In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for English and Chinese academic paper PDFs. The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model. The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies. We also provide a corresponding bilingual benchmark named PaperPDF.
We train PDF-WuKong on PaperPDF dataset, complemented by general-domain document question-answering datasets. Experimental results substantiate the effectiveness and efficiency of our approach to the task of long multimodal PDF understanding. PDF-WuKong significantly outperforms potential open-source models that may be applied to this task. It also surpasses some proprietary products for document understanding on our proposed PaperPDF benchmark. As the number of document pages increases, its accuracy and efficiency will not decrease significantly. It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE [17]. Besides, for the recent benchmark MM-NIAH [18] of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters. Our model achieves the best performance on multimodal content with a context length of 64K.
The main contributions of this paper are as follows:
  • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient PDF question answering.
  • We propose a bilingual PDF multimodal question answering dataset (PaperPDF) with $1.1M$ QA pairs for training and $10k$ QA pairs for evaluation.
  • Our model significantly surpasses existing open-source models and proprietary products (by an average of $8.6\%$ on F1) on long multimodal PDF understanding.
Input modality Type Number of tokens Models
Plain text Long-context Linear increase LongLoRA [4], LongLLaMA [5], YaRN [3]
RAG w/o Linear increase Graph RAG [9], DISC-LawLLM [6], RAPTOR [19]
Pure vision Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], XC2-4KHD [24], InternVL-V1.5 [11]
Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [25], CREAM [26]
Text and images Unlimited-page w/o Linear increase PDF-WuKong (Ours)
Table 1: Comparison of various models for processing multi-page long documents.
2 Related Works
2.1 Document Understanding Datasets
Early datasets focused on NLP tasks like summarization [27] and QA [28] of plain text, while visual document datasets targeted text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34]. Recent multimodal document QA datasets include DocVQA [35] and OCRVQA [36] for single-page documents, ChartQA [37] and ChartX [38] focusing on visual reasoning in charts. Datasets like ArXivQA [39] and InfoVQA [40] enhance MLLMs’ abilities on academic and infographic documents. However, these datasets are limited to single-page tasks, and current MLLMs [41, 11] perform well on them.
Multi-page QA datasets like MP-DocVQA [14], DUDE [17], DocGenome [42], and MM-NIAH [18] require understanding content relationships via multi-hop reasoning. Yet, answers in these datasets lack evidence and reliable interpretability, especially for questions needing multiple pieces of evidence from long documents.
2.2 Document Understanding Methods
Existing methods focus on plain text or limited document images. Text-based approaches convert documents into plain text using OCR and then employ long-context mechanisms like sparse attention [4], memory networks [5], or position interpolation [3]. Retrieval-augmented generation methods [6, 19, 9] also handle long texts effectively. These approaches struggle with fine-grained visual understanding.
Another solution, visual document understanding, treats each page as an image. MLLMs like UniDoc [20], mPLUG-DocOwl [21], and Vary [12] perform OCR-free understanding. Models such as UReader [22] and TextMonkey [10] divide high-resolution pages into patches. InternLM-XC2-4KHD [24] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration. However, reliance on high resolution increases token counts and isn’t scalable to multi-page documents.
For multi-page documents, models like Hi-VT5 [14], GRAM [16], Fox [13], CREAM [26] and mPLUG-DocOwl2 [25] encode pages separately and perform page-level interactions. More pages generate more visual tokens, increasing resource consumption and inefficiency for longer documents. Thus, we propose parsing documents into interleaved text and images, followed by sparse sampling in an end-to-end manner. Tab. 1 summarizes these methods.
3 Methodology
Figure 2: The overall structure of PDF-WuKong consists of a document parser, a sparse sampler and a large language model.
3.1 Overview
Our pipeline consists of three components: a document parser, a sparse sampler, and a large language model, as shown in Fig. 2. The document parsing stage converts input PDFs into machine-readable content with interleaved text and images. The sparse sampler then encodes and caches embeddings for text blocks and images separately. When receiving a user query, it retrieves the most relevant content through similarity matching. Finally, the query and sampled tokens are fed into the LLM for answer generation. The detailed procedure is outlined in Algorithm 1 in the appendix.
3.2 Document Parsing
Given a PDF document $D$, the goal of document parsing is to convert it into some machine-readable text blocks $\{T_{1},T_{2},\dots,T_{n}\}$ and diagrams $\{I_{1},I_{2},\dots,I_{m}\}$ according to the reading order and layout structure. By default, text blocks are organized into paragraphs, and all figures and tables are saved as images. These text and images are finally reorganized into an XML file in reading order. This process can be completed using existing open-source PDF parsing tools. During inference, we directly input the parsed structured full data into the subsequent stage of PDF-WuKong.
3.3 Sparse Sampling
For a lengthy multi-page document, if it is directly input into the LLM, there will be two problems. The first is the problem of computing efficiency. The consumption of computing resources will increase dramatically. The second is the problem of inaccurate attention. Key information related to the user query is easily submerged by a large amount of irrelevant content. It is difficult for the model to accurately locate and extract important information in a huge token sequence. Therefore, sparse sampling is essential for efficiently handling lengthy multi-page documents by identifying and extracting the most relevant text chunks or diagrams based on their similarity to the user query.
During the training, for the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query. Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec. 4). We randomly select two text blocks and two images from the remaining text blocks and images as negative samples. Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$. An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
Given the embeddings of the user query $e_{q}$, the positive samples $E_{P}=\{e_{T_{P}},e_{I_{P}}\}$ and negative samples $E_{N}=\{e_{T_{N}},e_{I_{N}}\}$, we employ a contrastive learning approach to align the text and image features with the query. The goal is to enable the model to capture the document content that is most relevant to the query. The contrastive learning loss is:
$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}}{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}+\displaystyle\sum_{e_{j}\in E_{N}}e^{\frac{\text{sim}(e_{q},e_{j})}{\tau}}},$ (1)
where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples. $\tau$ is the temperature parameter that controls the scale of the similarity scores. $P$ is the number of positive samples. By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
During the inference, we pre-encode all text blocks and images and cache all candidate embeddings. When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings. Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query. Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences. Moreover, the multimodal embedding cache further optimizes inference time. The pseudocodes for training and inference are shown in the appendix.
3.4 Answer Generation
At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency. Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence.
Considering that MLLM needs to encode images first for multimodal understanding, we directly input the image tokens obtained from the sparse sampler into the LLM, to save one image encoding process. Thus, the sparse sampler shares the same vision encoder with the MLLM. They can be integrated and trained in an end-to-end manner.
During the training, we input the positive text $T_{P}$ and the positive image tokens $e_{I_{P}}$ into the LLM. Besides, the query and instruction are also input into the LLM. Then, we calculate the cross-entropy loss $\mathcal{L}_{\text{QA}}$ between the output answer $a$ and the ground truth. Finally, the total optimization objective is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}.$ (2)
PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and QA.
4 PaperPDF Dataset
Figure 3: The construction process of PaperPDF based on single evidence and multiple evidence.
4.1 Overview
Our motivation for creating PaperPDF is threefold: (1) In long document QA contexts, answers often derive from specific segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multimodal PDF QA dataset. Therefore, we introduce a method for automatically generating question-answer pairs from long documents and present PaperPDF, a dataset designed for both training and evaluation.
4.2 Dataset construction strategy
The PaperPDF dataset is constructed through a four-step process as shown in Fig. 3: document parsing, evidence extraction, QA generation, and data filtering. First, $100k$ PDFs are parsed to extract text and image chunks. Then we extract some evidence from each PDF according to our specific rules. Given the evidence and the generation prompt as the input, we use Gemini Pro [43] to generate questions and answers for the training set due to its free and rapid accessibility. GPT-4V [44] is used for the test set to ensure high evaluation quality. After obtaining the preliminary dataset, we conduct automatic filtering and remove abnormal data. For the test data, we further perform manual filtering from multiple aspects to ensure its quality. According to different evidence extraction rules, the triplets in PaperPDF can be divided into two categories.
Single-evidence Q-E-A triplets. The questions in these triplets can be answered based on a single text chunk or a diagram. Therefore, the evidence can be categorized into Text-only and Image-only. These triplets enable PDF-WuKong to initially acquire the capabilities of sparse sampling and long multimodal document understanding.
Multi-evidence Q-E-A triplets. These triplets require reasoning across multiple chunks, involving combinations of text and images. The types include Image-text (derived from a text chunk and associated images), Section (generated from all chunks in a section), and Cross-paragraph (involving related paragraphs across a document). For the third type, semantic summarizations for each paragraph are conducted first, followed by a selection of several related text chunks for QA generation. These multi-evidence triplets enhance our model’s multi-hop reasoning capability.
In total, we obtained $1.1M$ bilingual training data and $10k$ testing data. The dataset statistics are shown in Tab. 2. The appendix contains more data statistics.
Category Train (En/Zh) Test (En/Zh)
Single Text-only 249k/12k 2939/296
Image-only 21K/40k 212/1018
Multi Image-text 250k/53k 2566/2150
Section 499k/7k 255/394
Cross-paragraph 1.2k/0 118/0
Table 2: The statistics of PaperPDF in English and Chinese.
5 Experiments
5.1 Implementation Details
The LLM and the vision encoder are initialized from IXC2-VL-4KHD [24] and the maximum number of dynamic tiles is set as 16. The text encoder is initialized from BGE-M3 [45]. We train the model by leveraging several document datasets including PaperPDF, DocVQA [35], ChartQA [37], InfoVQA [40], MPDocVQA [46], and DUDE [17]. Before both training and testing, PaperPDF is parsed using Grobid [47] and MinerU [48], while the other datasets are processed following their default instructions. The training is conducted for one epoch using 128 Ascend 910B NPUs with a learning rate of 4e-5. We set the top 5 sampling results as input to the LLM.
For the model evaluations on the PaperPDF dataset, we use three objective metrics ANLS, F1, and Rouge to report the quantitative results on the full test set. To evaluate the semantic correctness of the answer, we introduce GPT-Acc to determine whether the output is correct. Considering the expensive cost of GPT-4 evaluation, we randomly selected two subsets with 50 English PDFs and 30 Chinese PDFs for GPT-Acc calculation. They contain 488 and 317 QA samples, respectively.
Model #Param English Chinese Token
ANLS F1 Rouge GPT-Acc ANLS F1 Rouge GPT-Acc
Plain Text Solution
IXC2-VL [49] 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
IXC2-VL-RAG [49] 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
InternVL2 [11] 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
InternVL2-RAG [11] 8B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
Pure Vision Solution
Hi-VT5 [46] 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
IXC2-VL [49] 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
InternVL2 [11] 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
Parsed Interleaved Text and Images
IXC2-VL [49] 8B 27.8 31.2 32.6 37.7 22.5 29.5 29.2 31.4 6217
InternVL2 [11] 8B 33.4 36.2 36.6 54.3 28.5 40.6 42.0 54.7 6220
PDF-WuKong (ours) 8.5B 41.9 43.5 40.9 77.5 40.9 47.8 48.6 57.8 2107
Table 3: Performance comparison with open-source models for long PDF understanding on PaperPDF. The best results are marked bold and the second results are underlined.
Model ANLS F1 Rouge GPT-Acc
Gemini pro [50] 26.6 29.0 29.8 67.9
Kimi [51] 28.5 33.6 31.1 74.7
ChatGLM [52] 31.2 35.4 32.0 73.5
Qwen [53] 36.0 40.3 35.5 78.1
PDF-WuKong (ours) 41.8 43.2 40.7 77.5
Table 4: Performance comparison with commercial products (tested in Sep 2024) for long PDF understanding. The results are tested on a subset of 50 English PDFs. Note: These are the products based on the models rather than the models themselves.
5.2 Long PDF Understanding
To assess the effectiveness of our model in understanding long PDF documents, we conduct comprehensive experiments comparing it with both open-source models and commercial products on the PaperPDF dataset. Due to limited capabilities of traditional document understanding models  [46, 16, 54, 55, 56] and the huge resource costs in advanced MLLMs  [49, 24, 11], achieving deep understanding of lengthy PDF documents remains a highly challenging task. To handle this task, we explore three approaches to input the PDF documents into these MLLMs: pure text content, page images, and parsed interleaved text-image content. For some baselines with plain text input modality, we also report results based on retrieval enhancement techniques  [45]. The experimental results are reported in Tab. 3 and Tab. 4.
Several key conclusions can be drawn from the results. Firstly, inputting parsed interleaved text-image content generally outperforms multiple page images. The main reason is the limited input resolution and number of tokens that the model can accept. Secondly, when handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content. It is obvious that diagram information in the PDF plays a crucial role for document comprehension. This also indicates that PaperPDF places rigorous requirements on the visual information within documents. Additionally, we observe that for the InternVL2 model, the approach of parsing interleaved text-image content as input outperforms InternVL2-RAG. However, the opposite conclusion is drawn for IXC2-VL. We hypothesize that this discrepancy may be due to the max length setting in IXC2-VL, which could cause the model to overlook some critical information. Finally, benefiting from the inclusion of the spare sampler, our proposed PDF-WuKong model not only surpasses the existing state-of-the-art open-source model InternVL2 by approximately 7% on both the Chinese and English subsets, but also demonstrates competitive performance comparable to proprietary products. Moreover, due to the integration of the sparse sampler, PDF-WuKong maintains efficiency in inference token cost.
Model # param ANLS F1 ROUGE
Qwen-VL [57] 9.6B 26.4 19.6 18.3
Monkey [58] 9.8B 30.0 24.4 22.3
mPLUG-Owl2 [25] 8.2B 19.5 20.3 22.7
Emu2-Chat [59] 37B 26.0 24.4 23.4
MiniCPM-2.5 [60] 8.5B 31.8 28.2 24.8
IXC2-VL [49] 8B 23.4 20.8 21.3
IXC2-4KHD [24] 8B 24.5 20.0 18.0
CogVLM2 [61] 17B 24.8 27.4 26.3
[…]
Table 5: Performance comparison with other DocVLMs for PDF multimodal understanding on the Single-Evidence Subset. $\dagger$ denotes the page image input, aligning with other models in the table; * indicates that our model utilizes parsed content as input.
To compare with more open-source document MLLMs, considering that most of these models can only handle single-page documents, we construct a subset of the PaperPDF benchmark, only containing test samples with single evidence. Therefore, we provided all models with only one page containing the evidence as their input. The pages are input in the form of images. Our PDF-WuKong can accept input in two formats. One is the page image and another is the parsed page. As shown in Tab. 5, our model’s capability on this subset is significantly better than other document models. Moreover, the document parsing-based paradigm used in our approach is superior to the purely visual paradigm.
5.3 Document-oriented VQA
To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks and compare PDF-WuKong with other representative models. First, we evaluate the performance of PDF-WuKong on single-page document datasets [35, 37, 40]. As shown in Tab. 6, our model achieves comparable performance on single-page document understanding. This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question-answering tasks.
In addition, we assess the performance of traditional specialized models and MLLMs on two existing multi-page document QA datasets. The results shown in Tab. 7 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 [25]. Notably, on complex multi-page document datasets like DUDE [17], PDF-WuKong outperforms GPT-4V [44]. This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
Furthermore, we conduct zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18]. As shown in Tab. 8, our model uses the fewest parameters yet achieves the second-best performance. Although InternVL-V1-5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model. Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, while that of other models significantly decreases. At a context length of 64K, PDF-WuKong even achieves the best performance, demonstrating its robustness in handling long-context multimodal inputs.
Model DocVQA ChartQA InfoVQA
Qwen-VL [57] 65.1 65.7 35.4
Monkey [58] 66.5 65.1 36.1
Text-Monkey [10] 73.0 66.9 28.6
MiniCPM-V-2.5 [60] 84.8 - -
Vary-base [12] 76.3 66.1 -
TextHawk [62] 76.4 66.6 50.6
IXC2-4KHD-16 [24] 84.9 80.1 60.8
DocOwl 2 [25] 80.7 70.0 46.4
CREAM [26] 79.4 - 53.6
PDF-WuKong (ours) 85.1 80.0 61.3
Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks.
Model MP-DocVQA DUDE
LayoutLMv3 [54] 55.1 20.3
Longformer [55] 55.1 27.1
BigBird [56] 58.5 26.3
Hi-VT5 [46] 61.8 35.7
DocFormerv2 [63] 76.4 48.4
GRAM [16] 83.0 53.4
GPT-4V (2024-06) [44] - 53.9
Idefics3-8B [64] 67.2 38.7
DocOwl2 [25] 69.4 46.7
CREAM [26] 65.3 52.5
PDF-WuKong (ours) 76.9 56.1
Table 7: Performance comparison with other DocVLMs for multi-page document understanding.
Model #param Overall 1K 4K 16K 64K
Emu2-Chat [59] 37B 8.8 38.9 18.2 0.0 0.0
VILA1.0-13b [65] 13B 15.7 41.9 33.2 8.6 0.1
llava-v1.6-13b [23] 13B 16.9 43.7 34.9 13.6 0.0
llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
InternVL1.5-RAG [11] 45B 46.1 59.5 50.1 44.9 39.3
PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
Table 8: Performance comparison with other DocVLMs on MM-NIAH. The evaluation approach aligns with the benchmark.
5.4 Ablation Study
To comprehensively evaluate the effectiveness of our contributions, we conduct ablation studies focusing on the impact of the sparse sampler, the dataset, the document length, and sampling strategies. These experiments are based on English PaperPDF for training and evaluation to avoid interference from other factors.
Sparse sampler
To assess the effectiveness of the sparse sampler, we compared models trained with and without it. Without the sparse sampler, the MLLM struggled to process long documents with interleaved text and images, resulting in poor performance due to the large amount of irrelevant information. Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab. 9, by efficiently selecting the most relevant content for each query. Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately. This indicates that our end-to-end optimization of multimodal representation and question answering can further promote the document understanding ability of MLLM.
Sparse Sampler End-to-End ANLS F1 ROUGE
11.1 5.1 5.0
40.3 42.3 39.8
42.6 43.6 40.2
Table 9: Ablation study on the impact of sparse sampler
Dataset
We retrain PDF-WuKong on various subsets of our English PaperPDF, and the results are shown in Tab. 10. Increasing the amount of training data leads to consistent improvements in the model’s accuracy, proving that our dataset follows scaling laws. In addition, we verify the effectiveness of our two data construction methods. Under the same data scale, multi-evidence data can enhance the model’s complex reasoning ability.
Dataset ANLS F1 ROUGE
100 k 38.7 40.1 37.5
500 k 41.6 43.5 40.8
1 M 42.6 43.6 40.2
Single (200k) 39.2 40.4 38.1
Multi (100k) + Single (100k) 40.0 41.6 38.9
Table 10: Ablation study on dataset setting
Document length
To understand the impact of document length on model performance and efficiency, we divide the test set into subsets based on the number of pages per document. Results in Fig. 4 demonstrate that our model’s performance and token counts remain relatively stable across documents of varying lengths. This stability indicates that the sparse sampler effectively reduces the input size to a reasonable level, regardless of the original document length. In contrast, the baseline MLLM without the sparse sampler is unable to handle long documents effectively. Its performance deteriorates significantly as the document length increases. The number of tokens also increased dramatically, resulting in huge resource consumption. These findings highlight the robustness of our model in processing long documents without sacrificing accuracy or incurring extra computational costs.
Figure 4: Ablation study of different document length
Sampling strategy
We explore the impact of different numbers of text blocks or diagrams selected by the sparse sampler. As shown in Tab. 11, setting a small top $k$ can lead to missing crucial information needed for accurately answering queries, thus reducing performance. Conversely, a larger top $k$ introduces redundant information and increases computational costs without significantly enhancing accuracy. To strike a balance between performance and resource efficiency, we use the top 5 as the default setting.
Sampling chunks ANLS F1 ROUGE Tokens
Top 1 39.20 38.28 35.26 1186
Top 3 42.09 43.06 39.69 1452
Top 5 42.59 43.63 40.19 1789
Top 10 43.01 44.22 40.67 2386
Top 15 43.19 44.57 42.08 2704
Top 20 43.42 45.02 42.30 3364
Table 11: Ablation study of different sampling strategy
6 Conclusion
We have presented PDF-WuKong, a novel MLLM that effectively addresses the challenges of understanding long PDF documents containing interleaved text and images. By introducing an end-to-end sparse sampling mechanism, our model efficiently extracts the most relevant paragraphs and diagrams in response to user queries, significantly reducing input token size and making the process independent of document length. We also constructed PaperPDF, a bilingual dataset with $1.1M$ question-answer pairs for training and $10k$ pairs for evaluation, specifically tailored for academic PDFs. Experimental results demonstrate that PDF-WuKong not only outperforms existing open-source models but also surpasses proprietary products by an average of 8.6% in F1 score on the long multimodal PDF understanding task. Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
Appendix A Algorithm
Algorithm 1 shows the detailed inference process of PDF-WuKong. The training pipeline is shown in Algorithm 2. Our PDF-WuKong can achieve efficient and accurate understanding of long PDFs with end-to-end sparse sampling.
Algorithm 1 Inference pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$
2:Output: Generated answer $a$
3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
4:Stage 1: Document Parsing
5:Parse the input document $D$ into text blocks and images:
$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
6:Stage 2: Sparse Sampling
7:Encode all text blocks and images and cache all candidate vector embeddings:
$E_{T}=\{e_{T_{1}},e_{T_{2}},\dots,e_{T_{n}}\}\leftarrow\text{En\_T}(\{T_{1},T_{2},\dots,T_{n}\}),$
$E_{I}=\{e_{I_{1}},e_{I_{2}},\dots,e_{I_{m}}\}\leftarrow\text{En\_I}(\{I_{1},I_{2},\dots,I_{m}\})$
8:Encode the user query $q$:
$e_{q}\leftarrow\text{En\_T}(q)$
9:Calculate the similarity between query embedding $e_{q}$ and cached text/image embeddings $\{E_{T},E_{I}\}$:
$S_{T}=\{\text{Sim}(e_{q},e_{T_{i}})\mid i=1,2,\dots,n\},$
$S_{I}=\{\text{Sim}(e_{q},e_{I_{j}})\mid j=1,2,\dots,m\}$
10:Select the top-$k$ relevant text blocks and images:
$(T,I)_{top}\leftarrow\text{TopK}(S_{T},S_{I},k)$
11:Stage 3: Answer Generation
12:Input the query $q$ and the selected tokens into the LLM:
$a\leftarrow\text{LLM}(q,(T,E_{I})_{top})$
13:Return the generated answer $a$.
Appendix B Visualization Results
B.1 Qualitative Comparison with Other Products
The qualitative comparison between PDF-WuKong and other proprietary products on the long PDF understanding task is presented in Fig. 5. PDF-WuKong achieves significant performance advantages over others such as Qwen [66], ChatGLM [67], Kimi [68], and Gemini-pro [69]. The success stems from its ability to effectively sample key evidence relevant to the question and its strong multimodal understanding capabilities. This observation further validates the effectiveness of our method in enhancing the long PDF understanding capabilities of MLLMs.
Algorithm 2 Training pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$, ground truth answer $gt$
2:Output: Final loss function $\mathcal{L}_{\text{total}}$
3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
4:Stage 1: Data Preparing
5:Text blocks and images:
$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
6:Stage 2: Multimodal encoding
7:Encode the user query, positive and negative samples:
$e_{q}\leftarrow\text{En\_T}(q)$
$E_{T}=\{e_{T_{P}},e_{T_{N}}\}\leftarrow\text{En\_T}(\{T_{P},T_{N}\}),$
$E_{I}=\{e_{I_{P}},e_{I_{N}}\}\leftarrow\text{En\_I}(\{I_{P},I_{N}\}),$
8:Calculate the contrastive learning loss:
$\mathcal{L}_{\text{rep}}(e_{q},\{e_{T_{P}},e_{I_{P}}\},\{e_{T_{N}},e_{I_{N}}\})$
9:Stage 3: Output prediction of MLLM
10:Input the query, the positive text, and the positive image tokens from the shared image encoder En_I:
$a\leftarrow\text{LLM}(q,T_{P},e_{I_{P}})$
11:Calculate the cross-entropy loss:
$\mathcal{L}_{\text{QA}}(a,gt)$
12:Stage 4: Optimize model in an end-to-end manner
13:Update model parameters according to the joint loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}$
14:Return the final loss $\mathcal{L}_{\text{total}}$.
Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products. The red box indicates the evidence that the correct answer depends on.
B.2 Visualization Results on Chinese Documents
In addition to the outstanding performance on English documents, PDF-WuKong also demonstrates remarkable capabilities on Chinese documents. Fig. 6 shows its qualitative performance on the Chinese long PDF document understanding task. It effectively samples relevant evidence within the PDF and demonstrates the ability to derive accurate answers based on the sampled evidence.
Appendix C More Dataset Details
The Q-E-A triplets in the PaperPDF dataset are derived from approximately 60k English and 10k Chinese documents, encompassing nearly 70 disciplines such as computer science, engineering, and materials science.
To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis. The key statistics of the source documents in PaperPDF are presented in Tab. 12. Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents but also features more pages and OCR tokens. More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images. Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task.
For the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence, Fig. 7 - Fig. 11 show their prompt engineering and a corresponding example, respectively.
Figure 6: Examples of PDF-WuKong on Chinese documents. The red box indicates the evidence that the correct answer depends on.
MPdataset Document Language Page Image Token
DUDE [17] 5k En 6 - 1831
MP-Docvqa [46] 6k En 8 - 2026
PaperPDF (En) 60k En 25 12 11371
PaperPDF (Zh) 10k Zh 11 7 3413
PaperPDF 70k En+Zh 23 11 10234
Table 12: Statistics of the source documents, where Page, Image, and Token represent the average number of pages per document, the average number of images in the document, and the average number of OCR tokens per document, respectively.
Text-only Question Generation Prompt Task Definition: Create 2 academic questions from a given research paper paragraph. Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions. Focus on the paragraph,not the entire paper. If the paragraph lacks valid information,return ‘quit’. You should use English. Develop 2 questions that: Are no more than 30 words. Incorporate knowledge from the paragraph. Are answered by text instead of one of the multiple choices. Are elicit detailed responses supported by the text. Expected Output: (Return ‘quit’ directly if the paragraph lacks valid information.) [Q1]: question1 here [Q2]: question2 here
Text-only Answer Generation Prompt Task Definition: Answer a question based on the material given. Requirements: The answers should: Be comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than 20 words. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model? Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig. 5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig. 5b). Answer 1: It can lead to an unfair advantage for the model. Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset. If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
Figure 7: Text-only Q-E-A triplets generation prompt and data example.
Image-only Question Generation Prompt Task Definition: Formulate 2 academic questions based on a provided figure or table from a research paper. Requirements: The questions must directly reference and integrate information presented in the image and its caption, ensuring a cohesive understanding of the content depicted. You should use English. Develop 2 questions that are: No more than 30 words. Specific to the unique data or details visible in the figures/tables and are answerable only based on the material without inferring or speculating on details not explicitly explained by the figures/tables. Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’. Expected Output: [Q1]:[Q2]:
Image-only Answer Generation Prompt Task Definition: Answer a question based on an image and its caption from a research paper. Requirements: The answers should: Always use English. Not infer or speculate on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words. [thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
Image-only Data Example Query: Based on Figure 4, which fuzzing technique consistently leads in coverage across all benchmarks over the 24-hour period? Figure: Figure 4Refer to caption Caption: Figure 4: Number of benchmarks on which each technique has the lead in coverage at each hour. A benchmark is counted for multiple techniques if two techniques are tied for the lead. Answer 1: FairFuzz consistently leads in coverage across all benchmarks over the 24-hour period in Figure 4. Answer 2: FairFuzz, highest coverage benchmark count over 24 hours.
Figure 8: Image-only Q-E-A triplets generation prompt and data example.
Text-image Question Generation Prompt Task Definition: Formulate 2 academic questions based on a provided paragraph, a figure or table from a research paper. Requirements: The questions should not rely on the accompanying text but only the figure / table. However, you may use the provided text to understand the figure / table. You should use English. Develop 2 questions that are: No more than 30 words. Based on a provided figure of a research paper, without relying on accompanying text for the questions. However, you may use the provided text to understand the figure. Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’. Expected Output: [Q1]:[Q2]:
Text-image Answer Generation Prompt Task Definition: Answer a question based on the material given. Requirements: The answers should be: Always using English. Comprehensive and cover all relevant aspects, as presented in the provided text and figure. Accurately reflect the information and insights offered by the research paper. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words. [thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1Refer to caption Caption: Experimental results of ablation analysis, … Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Figure 9: Text-image Q-E-A triplets generation prompt and data example.
Section Question Generation Prompt Task Definition: Formulate 2 academic questions based on a section from a research paper. Requirements: Carefully read and comprehend the entire provided section of the research paper to ensure a thorough understanding of its content, including key points, findings, methodologies, and conclusions. You should Always use English. Develop 2 questions that are: No more than 30 words. Requiring an integration of information from all paragraphs and figures/tables in the section. Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’. Not based on common knowledge or assumptions not supported by the figures and tables. Expected Output: [Q1]:[Q2]:
Section Answer Generation Prompt Task Definition: Answer a question based on the material given. Requirements: The answers should be: Always using English. Not inferring or speculating on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words. [thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
Section Data Example Query: How does the qualitative evaluation of extractive summarizers using word clouds elucidate the differences in content focus between the original documents and the summaries? Text:
Here we use word cloud representations to give an intuitive interpretation of the content in the generated extractive summarizers…
Figure 3 shows a word cloud made by the aggregation of all the summaries generated by the PKUSUMSUM-Centroid method… The images clearly show a contrast of content… Figure: Figure 3Refer to caption Caption: Figure 3: The word clouds representing summaries generated by PKUSUMSUM-Centroid method (left) and original documents without the content of those summaries (right). Answer 1: Visual contrast in word frequency highlights content focus differences. Answer 2: Word clouds highlight the prominent themes in summaries versus original texts by displaying relative word frequencies visually.
Figure 10: Section Q-E-A triplets generation prompt and data example.
Cross-paragraph Question Generation Prompt Task Definition: Based on the selected paragraph from a research paper that share a thematic or conceptual connection, formulate an insightful, open-ended question. This question should reflect the shared themes or concepts of your selections and relate to the broader context of the research paper. Requirements: Ascertain the underlying connection among the paragraphs and the figures/tables(if provided). Subsequently, craft an insightful, open-ended question that encapsulates the identified themes or connections, aiming to foster analytical thinking and in-depth discussion on the subject matter of the paper. Expected Output: [Q]: [Your generated question based on the shared themes or information]
Cross-paragraph Answer Generation Prompt Task Definition: Given some selected paragraphs from a research paper, and ensuring that these paragraphs share a certain level of association, you are to answer a question that is related to the content of these selected paragraphs. Requirements: Craft the 2 answers for the question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Expected Output: [A1]: [Insert a concise answer here, no more than 20 words.] [A2]: [Insert a detailed answer here,including a detailed “thought chain” or reasoning process.]
Cross-paragraph Data Example Query: How can we leverage the proposed EDO approach to optimize the selection of datasets for specific algorithms, thereby enhancing the overall performance and validity of the algorithms? Text:
Paragraph 7:
Figure 1: On the right: the current path for selecting some algorithm(s), …
Paragraph 17: Section 2 describes the structure of the proposed method including its parameters and operators. Paragraph 20: In this section, the details of an algorithm that generates data for which a given function, … Figure: Figure 1Refer to caption Caption: Sample number of rows and columns 2. Sample columns and fill/trim values as needed. Answer 1: The EDO approach identifies optimal datasets, improving algorithm performance and validity through tailored data selection. Answer 2: We can systematically evaluate and select datasets that align closely with the specific requirements of the algorithms in use. This method analyzes various dataset characteristics, ensuring that the chosen data not only matches the algorithm’s operational parameters but also enhances its predictive accuracy.
Figure 11: Cross-paragraph Q-E-A triplets generation prompt and data example.
Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model? Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig. 5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig. 5b). Answer 1: It can lead to an unfair advantage for the model. Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset. If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1[Uncaptioned image] Caption: Experimental results of ablation analysis, different training strategies of our method and comparison with other approaches. Note: the -means the results were not reported by that method. Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Appendix D Limitation
For our proposed dataset, current documents are mainly limited to academic papers, so the layout format and subject matter of the documents are relatively simple. We will expand the dataset with more diverse documents. Besides, our model is not specifically designed for some global queries, which will become a key research problem in the future.